OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 LINE 1: y = _____x + _____ LINE 2: y= _____x + _____

OpenStudy (anonymous):

this one i have now clue at all plz help me understand this one @livias.random

OpenStudy (anonymous):

So the form they are looking for is y=mx+c where m is the slope (change in y/change in x)

OpenStudy (anonymous):

ok i am writing this down for sure

OpenStudy (anonymous):

Then you work out c by (y-b) = m (x-a)

OpenStudy (anonymous):

This is how I've been solving all the other ones

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i will try to solve it i will show it to u when done

OpenStudy (anonymous):

Sure :) If I'm slow to respond though, don't worry

OpenStudy (anonymous):

ok i think i have it is it: y= -2x + 1 y= -2x + 5 ?????????????????????????????????????????????????? hope i am right

OpenStudy (anonymous):

@livias.random i have it i hope i am right can u check?

OpenStudy (anonymous):

m = (7-5)/(-3-(-2)) m = 2/-1 m = -2 (y-7) = 2(x-(-3)) y-7 = 2x + 6 y = 2x + 13 - m = (5-1)/(0-2) m = 4/-2 m = -2 (y-5) = -2(x-0) y-5 = -2x y = -2x + 5 I got slightly different for the first but I rushed it a bit so I'll double check

OpenStudy (anonymous):

aw i got the second one right but not the first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not bad :)

OpenStudy (anonymous):

it is all right both of them?

OpenStudy (anonymous):

i have one more i will solve before i post though :) here is medal

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

np :)